OpenStudy (anonymous):

Evaluate the expression: [(8 × 4) − 10]2 + 9

OpenStudy (anonymous):

need help

OpenStudy (souvik):

8*4=?

OpenStudy (anonymous):

32

OpenStudy (souvik):

[(8*4)-10]=?

OpenStudy (anonymous):

22

OpenStudy (souvik):

good !now [(8*4)-10]*2=?

OpenStudy (anonymous):

44

OpenStudy (souvik):

now [(8 × 4) − 10]2 + 9 =?

OpenStudy (anonymous):

53

OpenStudy (karatechopper):

PEMDAS! [(8 × 4) − 10]2 + 9 Parenthesis [(8 × 4) − 10]2 + 9 [(32)-10]2+9 Finish out the bracket [22]2+9 Multiplication/Division 22(2)+9 44+9 Addition/Subtraction 44+9=53

OpenStudy (souvik):

yeah!right....:)

OpenStudy (anonymous):

for some reason my thing dont give that answer- heres my answers A. −59 B. 493 C. 933 D. 32

OpenStudy (souvik):

there is nothing wrong in what u did... make sure that the u type here is right..

OpenStudy (anonymous):

i typed in the question right

OpenStudy (karatechopper):

Completely sure?

OpenStudy (anonymous):

yes double checked

OpenStudy (karatechopper):

hmph, then i don't know what to do. :/

OpenStudy (anonymous):

i think i figured it out but thanx anyways :)

OpenStudy (souvik):

:)

OpenStudy (karatechopper):

:D